OpenStudy (anonymous):

What is the solution to the equation 1 over the square root of 8 = 4^(m + 3) ?

OpenStudy (tkhunny):

\(\dfrac{1}{\sqrt{8}} = 4^{m+3}\) Well, \(\dfrac{1}{\sqrt{8}} = 2^{-2/3}\) and \(4^{m+3} = 2^{2m+6}\). Anything there?